### v1

# footnotetext: *Equal Contributionfootnotetext: Corresponding authors

Abstract
Effective clinical decision-making depends on iterative, multimodal reasoning across diverse sources of evidence. The recent emergence of multimodal reasoning models has significantly transformed the landscape of solving complex tasks. Although such models have achieved notable success in mathematics and science, their application to medical domains remains underexplored. In this work, we propose MedE2, a two-stage post-training pipeline that elicits and then enhances multimodal reasoning for medical domains. In Stage-I, we fine-tune models using 2,000 text-only data samples containing precisely orchestrated reasoning demonstrations to elicit reasoning behaviors. In Stage-II, we further enhance the model’s reasoning capabilities using 1,500 rigorously curated multimodal medical cases, aligning model reasoning outputs with our proposed multimodal medical reasoning preference. Extensive experiments demonstrate the efficacy and reliability of MedE2 in improving the reasoning performance of medical multimodal models. Notably, models trained with MedE2 consistently outperform baselines across multiple medical multimodal benchmarks. Additional validation on larger models and under inference-time scaling further confirms the robustness and practical utility of our approach.
1 Introduction
Medicine is a multifaceted endeavor involving daily diagnostic and therapeutic activities performed by healthcare professionals. It requires clinicians to process vast amounts of information, including patient medical histories, physical examination findings, and laboratory test results, to diagnose conditions, formulate prognoses, and determine appropriate treatment plans. In many clinical settings, an iterative reasoning process that evaluates multiple possibilities with progressively accumulated clinical information is considered fundamental to medical practice adler2021next ; singh2022five ; croskerry2022advancing . Recent advancements in multimodal large language models (MLLMs), such as OpenAI-o-series openai2024reasoning and Gemini-2.5-Pro Google2025Gemini , have significantly advanced complex task performance. These models employ scaling inference time and emulate reflective cognitive processes, pushing capabilities to unprecedented levels. In light of these advancements, we explore the question: How can we effectively extend the strategy of multimodal reasoning to medical domains?
Figure 1: (a) General MLLMs consistently show superiority over models specifically designed for medical domains across multiple datasets, especially on complex and diverse tasks. (b) Samples that closely mirror real-world clinical scenarios are used to strengthen multimodal reasoning capabilities instead of samples focused primarily on pattern recognition or basic knowledge recall.
Current multimodal reasoning models demonstrating remarkable proficiency are built upon powerful multimodal foundation models bai2025qwen2 ; chen2024internvl that perform well across various general benchmarks yue2024mmmu ; yue2024mmmupro ; meng2025mm ; he2024olympiadbench ; lu2023mathvista . We begin by assessing the capabilities of different models within the medical domain. Such evaluations represent a natural starting point for probing a model’s medical knowledge and reasoning capacities. The results in Figure 1 indicate that current multimodal models specifically tailored to medical tasks chen2024huatuogpt ; li2023llava typically focus only on particular applications, such as visual question answering and medical report generation, and exhibit performance degradation in other medical tasks. In contrast, general multimodal large language models demonstrate superior results. Therefore, we adopt general MLLMs as a foundation for developing specialized medical reasoning models.
Enhancing a model’s capacity for rigorous reasoning depends far more on the quality and domain-specific relevance of its training corpus than on the sheer volume of samples  deepseekai2025deepseekr1incentivizingreasoningcapability . General models predominantly use training samples from mathematical and scientific domains. In contrast, while medical practice encompasses numerous scenarios requiring reasoning, the available data is limited. Unlike previous studies that rely on automatically generated simple question-answer pairs, we implement a rigorous data construction process involving meticulous data collection, cleaning, and expert review. Specifically, we exclude samples that are solvable solely through inherent knowledge or pattern recognition, as shown in Figure 1. Consequently, our carefully curated dataset comprises 2K textual questions sourced from Medical Licensing Examinations jin2021disease and 1.5K multimodal medical questions integrating clinical histories, physical examinations, investigations, and procedures across 12 imaging modalities, including radiology, pathology and optical coherence tomography.
With the dataset in place, we proceed to develop the models’ reasoning capabilities specifically for clinical scenarios. Previous work meng2025mm ; deepseekai2025deepseekr1incentivizingreasoningcapability ; su2025gmai ; li2025limr ; ye2025limo usually employs rule-based reinforcement learning shao2024deepseekmath to advance models’ reasoning capabilities. However, challenges such as normalization of specialized medical terms (e.g., anatomical terminology) and hierarchical relationships (e.g., organ-system classifications) in medical domains make it difficult to verify responses solely through predefined rules. Moreover, methods that utilize neural reward models or reinforcement learning via multiple-choice questions are susceptible to reward hacking denison2024sycophancy . Building upon these, we introduce a novel training pipeline, named MedE2. Our pipeline involves two stages to progressively elicit and enhance the clinical cognitive reasoning capability of the models. In the first stage, we conduct supervised fine-tuning on text-only data, where each sample includes carefully designed reasoning demonstrations to show the behavior that how to utilize existing knowledge to solve complex tasks. Subsequently, we enhance the quality of the generated reasoning process to align with desired patterns. During this stage, we formulate Multimodal Medical Reasoning Preference (MMRP) and utilize Gemini-2.5-Pro to perform rejection sampling on our curated dataset, while employing Direct Preference Optimization (DPO) ziegler2019fine ; ouyang2022training to calibrate and refine the model’s capabilities.
Extensive experiments are conducted to validate the effectiveness of our proposed pipeline, MedE2. Despite being trained only on text-based reasoning data during Stage-I, the models demonstrate significant performance improvements across multiple medical benchmarks, achieving gains of at least 4.45% on MedXpertQA-MM zuo2025medxpertqa and 6.67% on MMMU-Pro-Health yue2024mmmupro . These results highlight an innovative method for eliciting reasoning behaviors in multimodal medical tasks. Further enhancement in Stage-II enables models to exhibit superior reasoning capabilities, reaching performance competitive with several larger-scale models. We also show that improvements from MedE2 generalize well to models with larger parameters, underscoring the scalability of our approach. Moreover, the consistent improvements observed under inference-time scaling further demonstrate the robustness of our post-training recipe.
In summary, our main contributions are as follows:
  • We present MedE2, a two-stage post-training pipeline designed to enhance multimodal reasoning in medical scenarios. The pipeline comprises two critical stages: (1) eliciting reasoning behavior by leveraging strategically developed text-only datasets, and (2) refining the quality of reasoning by incorporating meticulously curated multimodal data.
  • We construct a high-quality dataset of approximately 3.5K samples, comprising both text-based questions with reasoning chains and multimodal questions involving clinical information, thereby establishing a reliable corpus for bootstrapping medical reasoning.
  • We perform comprehensive experiments across diverse benchmarks to validate our post-training pipeline. Experimental results consistently demonstrate that MedE2 significantly enhances model performance and achieves strong generalization across models of varying sizes, while also being robust to inference-time scaling.
2 Related Work
2.1 Medical Multimodal Large Language Models
The development of multimodal large language models has significantly advanced the field of medicine. Current medical multimodal large language models are primarily based on general multimodal models and are further trained on specialized medical datasets. This approach led to the emergence of numerous medical multimodal large models. For example, Med-Flamingo moor2023med builds upon OpenFlamingo-9B awadalla2023openflamingo and continues pretraining with medical image-text data, while Med-PaLM tu2024towards constructs the MultiMedBench dataset and fine-tunes based on PaLM-E driess2023palm . Recent studies continuously use similar strategies and refined training methods, resulting in significant progress, such as LLaVA-Med li2023llava , BioMedGPT zhang2024generalist , MedTrinity-25M xie2024medtrinity , Qilin-Med-VL liu2023qilin , and Med-Gemini saab2024capabilities . These strategies have demonstrated good application results in various medical scenarios, including medical dialogue singhal2025toward , clinical decision support schubert2023evaluating , electronic health record analysis luo2022biogpt , and image report generation li2018hybrid . Given the successful application of reasoning capabilities in scientific tasks, their potential in the medical field is vast; however, effective training strategies to enhance the reasoning abilities of medical models are still lacking. Therefore, our study aims to explore how to enhance the reasoning abilities of multimodal large models in medical tasks.
2.2 Reasoning Models
Enhancing the reasoning abilities of models remains a key challenge. Early efforts evolved from few-shot prompting to structured paradigms like Chain-of-Thought wei2022chain and ReAct yao2023react , aiming to emulate human-like reasoning. Subsequent work recognized reasoning as an iterative trial, error, and refinement process. This led to the introduction of heuristic search and the process-reward model. Heuristic search, inspired by algorithms such as Monte Carlo Tree Search silver2017mastering , extended single reasoning chains into tree-structured thoughts yao2023tree ; shinn2023reflexion . In parallel, process-reward models framed reasoning as a Markov Decision Process, assigning stepwise rewards to guide the generation of more accurate reasoning chains lightman2023let ; jiao2024learning ; wang2023math . The release of OpenAI’s o1 openai2024reasoning catalyzed further progress. Journey Learning qin2024o1 explored strategies for o1-style slow thinking; o1-coder zhang2024o1 incorporated tree search in code tasks; and STILL-2 min2024imitate distilled long-form reasoning data to expand and refine solution paths. These advancements, culminating in DeepSeek’s results deepseekai2025deepseekr1incentivizingreasoningcapability , highlight the critical role of reinforcement learning in improving reasoning. Diverse reinforcement learning approaches to further bolster reasoning are now being actively explored and investigated wang2024enhancing ; wei2025skywork ; su2025gmai .
Figure 2: Overview of the two-stage post-training recipe MedE2. In Stage-I, text-only data containing reasoning demonstrations is employed to elicit initial reasoning behavior in both textual and multimodal scenarios. In Stage-II, Direct Preference Optimization is applied to multimodal data to further enhance reasoning quality, aligning the model’s outputs more closely with human preferences.
3 Pipeline
This section presents the core pipeline for adapting multimodal reasoning strategies to the medical domain, as illustrated in Figure 2. To enhance the model’s reasoning capabilities, we first curate a high-quality dataset comprising challenging text and multimodal question-answer pairs (Section 3.1). Based on our pilot studies, we select general MLLMs (e.g., QwenVL2.5 bai2025qwen2 and InternVL3.0 chen2024internvl ) as the base models due to their balanced performance across various medical tasks. Building upon these base models, we propose a two-stage training paradigm that progressively enhances the model’s ability to reason in complex clinical scenarios (Section 3.2 and Section 3.3).
3.1 Data Collection
We begin by constructing a large-scale question bank sourced from authoritative examinations, top-tier journals, and publicly available datasets. In total, we collect 54K samples, consisting of 37K textual-only and 17K multimodal questions. To ensure high quality and prevent data leakage, a rigorous filtering process was implemented in collaboration with AI experts. Specifically, we employ two models—GPT-4o as the baseline model and Gemini-2.5-Pro as the advanced model—for sample screening. Given that our goal is to explore the potential of multimodal reasoning strategies in solving complex medical tasks, we prioritize filtering out samples lacking sufficient reasoning complexity.
For textual questions, we initially remove samples that GPT-4o could solve directly, indicating they required no meaningful reasoning. The remaining questions are further evaluated by Gemini-2.5-Pro, which attempts each question 4 times. We retain only those samples that the model answers correctly at least once. For multimodal questions, we first assess whether the problem can be solved through mere pattern recognition by prompting GPT-4o to determine the question type. Subsequently, two AI experts conduct independent sampling attempts on the remaining questions. We preserved only those questions correctly answered by at least one expert across all attempts. Ultimately, this process resulted in a dataset consisting of 3.5K samples, including 2K textual and 1.5K multimodal questions.
3.2 Stage-I: Eliciting Reasoning Ability
To enable models to effectively reason in clinical problem-solving scenarios, it is essential to instruct them on utilizing their existing knowledge base to address complex reasoning tasks. Previous studies ye2025limo ; li2025limr have demonstrated that while supervised fine-tuning with reasoning-specific data can enhance reasoning task performance, reasoning capabilities can be elicited effectively with only a few illustrative examples. Indeed, since much of the relevant knowledge has already been encoded during pre-training, the construction of precisely orchestrated demonstrations of reasoning processes is more critical than merely increasing the volume of training data.
In this stage, we utilize textual-only reasoning demonstrations to elicit sophisticated reasoning capabilities, rather than relying on multimodal reasoning data. This choice is motivated by empirical findings: (1) supervised fine-tuning on multimodal datasets tends to encourage memorization rather than generalization, and (2) entangled multimodal reasoning data can impair the model’s original language reasoning abilities. In contrast, training exclusively on textual data not only enhances reasoning skills but also maintains general visual understanding, albeit with a slight trade-off.
Building upon prior efforts, we adopt a distillation-based method to generate high-quality reasoning demonstrations using our curated dataset. Specifically, we leverage state-of-the-art reasoning models, including o1 openai2024reasoning , DeepSeek R1 deepseekai2025deepseekr1incentivizingreasoningcapability , and QwQ-32B qwq32b , to produce diverse solutions. For open-source models, we directly use their generated reasoning processes. For the proprietary model, such as o1, whose intermediate reasoning steps are not readily accessible, we further prompt the model to expand its summarized outputs into complete reasoning processes. We conduct rigorous evaluations that combine rule-based filtering with human-assisted validation to ensure the quality of generated solutions. We design three criteria: (i) the correctness of the final answer, (ii) logical structure and organization, and (iii) the plausibility and coherence of the reasoning process. In addition, we select a subset of multimodal instances and prompt o1 to elaborate the relevant explanations into detailed reasoning processes, which are filtered using the same criteria and used for comparison.
3.3 Stage-II: Enhancing Reasoning Quality
Although supervised fine-tuning can elicit reasoning behavior in models, it is often accompanied by an increase in hallucinations. Prior work lv2024coarse ; akbar2024hallumeasure attributes this phenomenon to a mismatch between the generation paradigms during training and inference. Specifically, models are trained to predict the next token conditioned on preceding ground-truth tokens, while at inference time they rely on their own previously generated outputs. This discrepancy becomes particularly pronounced in tasks requiring long-form reasoning. As illustrated in Figure 2, even a correct answer can emerge from a flawed reasoning process involving terminology misuse, logical errors, or fabricated medical content, etc. To align model outputs with human preference, the mainstream method is to employ Reinforcement Learning from Human Feedback (RLHF) ziegler2019fine ; ouyang2022training , exemplified by Preference optimization (PO). In this work, we introduce Multimodal Medical Reasoning Preference (MMRP) and integrate it with Direct Preference Optimization (DPO) rafailov2023direct ; xu2024dpo . This combination effectively reduces hallucinations and generates reasoning processes that align more closely with user requirements.
Preliminary Considering a model as a policy $\pi_{\theta}(y|x)$ parameterized by $\theta$, RLHF aims at aligning the LLM $\pi_{\theta}$ with human preference. DPO is a representative algorithm of RLHF, serving as the preference loss to optimize the policy $\pi_{\theta}$ by enabling the model to learn the relative preference between chosen and rejected responses. DPO eliminates the requirement of training an explicit reward model based on the assumption of the Bradley-Terry model huang2004generalized and directly optimizes $\pi_{\theta}$:
$\mathcal{L}_{\mathrm{DPO}}\left(\pi_{\theta}\right)=-\mathbb{E}_{\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l}\right)\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{w}\mid\mathbf{x}\right)}{\pi_{\mathrm{ref}}\left(\mathbf{y}_{w}\mid\mathbf{x}\right)}-\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{l}\mid\mathbf{x}\right)}{\pi_{\mathrm{ref}}\left(\mathbf{y}_{l}\mid\mathbf{x}\right)}\right)\right]$
where $\mathbf{y}$ denotes the response of $\mathbf{x}$, and $\mathbf{y}_{w}$, $\mathbf{y}_{l}$ represent the “win” and “lose” items in preference pairs. $\sigma$ is the sigmoid function. $\pi_{ref}$ is the reference model used to regularize $\pi_{\theta}$ via Kullback–Leibler divergence, with $\beta$ controlling the strength of regularization. DPO directly assigns higher probabilities to preferred responses, aligning with human preferences, while bypassing the train-inference mismatch present in SFT.
Multimodal Medical Reasoning Preference The reasoning process is logically coherent and step-by-step valid. Contain appropriate analysis of relevant visual information. Avoid introducing hallucinated content not grounded in the input. Includes self-checking, verification, or reflection on uncertainties.
Preference Data Construction. Initially, we prompt the Stage-I model $M_{1}$ to generate multiple candidate reasoning processes for each sample in our multimodal training dataset. Given an image $I$ and a question $q$, we sample candidate reasoning processes $y$ from the distribution $M_{1}(y\mid q,I)$, repeating this process 8 times per sample to form a candidate set $Y$. Notably, each question $q$ undergoes meticulous refinement through collaboration between human experts and a model (GPT-4o), explicitly removing descriptions of image $I$ from $q$. This ensures that reasoning arises from actual image examination rather than from textual captions. Subsequently, we employ GeminiPro2.5 to evaluate these candidate reasoning processes according to five MMRP criteria: Answer Correctness, Logical Consistency, Image Analysis Involvement, Absence of Hallucinations, and Presence of Reflection. Reasoning processes meeting all criteria form the positive set ($Y_{p}$), whereas those failing any criteria constitute the negative set ($Y_{n}$). We then construct preference pairs by selecting a preferred response $y_{c}$ from $Y_{p}$ and contrasting it with a response $y_{r}$ from $Y_{n}$. Unlike static dataset construction, our MMRP methodology enables dynamic and flexible curation of multimodal preference datasets, effectively balancing specialization and generalization in reasoning tasks. The prompt used for data evaluation with Gemini-2.5-Pro can be found in Appendix A.
After applying DPO-based fine-tuning to the constructed preference dataset, the model demonstrates improved alignment with desired reasoning patterns and shows enhanced performance across diverse multimodal clinical scenarios.
4 Experiments
4.1 Experimental Setup
We select QwenVL2.5-7B and InternVL3.0-8B as the base models to evaluate the effectiveness of MedE2, resulting in two variants: (1) +Stage-I, which involves supervised fine-tuning on text-only data containing reasoning demonstrations; (2) +Stage-II, where the models from Stage-I are further enhanced using multimodal data and trained with DPO. We report the performance of GMAI-VL-R1 su2025gmai as a baseline method for comparison. For context, we report results from several leading proprietary models—GPT-4o, OpenAI-o1 openai2024reasoning , Gemini-2.5-Pro Google2025Gemini , and QvQ-Max teamqvq —to provide strong reference baselines. We also benchmark our approach against state-of-the-art open-source models that operate at substantially larger parameter scales bai2025qwen2 ; chen2024internvl .
All experiments are conducted on 8 NVIDIA H100 GPUs. LLaMA-Factory zheng2024llamafactory is used as the training framework, and vLLM kwon2023efficient serves as the backend during inference. In Stage-I, models are fine-tuned using LoRA hu2022lora with a learning rate of $1\times 10^{-4}$. Stage-II involves full-parameter fine-tuning with a learning rate of $1\times 10^{-6}$, optimized using DeepSpeed rasley2020deepspeed with the ZeRO-3 configuration. Both stages adopt a cosine learning rate decay schedule.
4.2 Evaluation Benchmarks
Our experiments involve two text-only tasks—MedQA jin2021disease and Medbullets chen2025benchmarking —as well as three multimodal benchmarks: MedXpertQA-MM zuo2025medxpertqa , MMMU-Health yue2024mmmu , and MMMU-Pro-Health yue2024mmmupro . Both MedQA and Medbullets are derived from questions used in the United States Medical Licensing Examination (USMLE). MedQA includes questions from both Step 1 and Step 2/3, while Medbullets focuses solely on Step 2/3, which are generally considered more challenging. For multimodal medical evaluation, MedXpertQA-MM serves as a highly challenging benchmark that assesses both medical knowledge and reasoning. MMMU-Health and MMMU-Pro-Health are healthcare-specific subsets of MMMU and MMMU-Pro, respectively.
4.3 Main Results
Text-only SFT Elicits Reasoning Behavior. As shown in Table 1, even a limited number of samples formatted with extended reasoning chains can effectively elicit the reasoning behavior, resulting in substantial performance gains. For example, QwenVL2.5-7B achieves improvements of +4.45% on MedXpertQA-MM and +6.67% on MMMU-Health. More impressively, the performance improvements from reasoning elicitation are even more pronounced for stronger base models. For example, InternVL3.0-8B achieves an 8% gain on MMMU-Health. This observation aligns with our hypothesis that training samples can serve as templates, illustrating how to utilize existing knowledge to solve complex reasoning tasks. Compared to GMAI-VL-R1 SFT, which relies on a 10K multimodal samples, our method achieves a 6 percentage point increase on MMMU-Health and a 3.37 percentage point gain on MMMU-Pro-Health only one-fifth of the data in a text-only format. This underscores that high-quality, task-targeted supervision is more critical than sheer data quantity for developing robust reasoning abilities.
Multimodal DPO Enhances Reasoning Quality. When Direct Preference Optimization (DPO) is applied after SFT in Stage-II, it builds upon the structured reasoning patterns established during the earlier stage and further enhances the model’s output quality. Compared to Group Relative Policy Optimization (GRPO)-based tuning shao2024deepseekmath (e.g., GMAI-VL-R1 RLT), our method achieves notably better results: 26.55% vs. 23.80% on MedXpertQA-MM, 66.00% vs. 57.33% on MMMU-Health, and 38.81% vs. 33.45% on MMMU-Pro-Health. A similar pattern is also observed in InternVL3.0-8B. These results suggest that merely forcing the model to reason over limited-solution-space tasks during training is suboptimal, often leading to hallucinations or incoherent reasoning. Instead, activating reasoning capabilities requires not only complex tasks but also training samples that engage the model in inference-time computation, or in other words, examples with precisely orchestrated solutions. Although the performance gain from DPO is slightly smaller than those from text-only SFT, we observe that DPO helps regulate the issue of “endless thinking” and encourages the model to “look before it thinks.” Qualitative analysis in Section 6 further shows that models trained with Stage-II generate more coherent and sophisticated reasoning processes.
Table 1: Results (%) of applying MedE2 to QwenVL2.5-7B and InternVL3.0-8B on multimodal medical benchmarks, along with state-of-the-art methods and other training strategies. $\Delta$ indicates the performance gain over the base model, and the best improvements are highlighted in bold.
Method MedXpertQA-MM MMMU-Health MMMU-Pro-Health
Reasoning Understanding Overall $\Delta$ Overall $\Delta$ Overall $\Delta$
Proprietary
GPT-4o 40.73 48.19 42.80 59.33 40.91
QvQ-Max 35.20 38.99 36.25 70.67 52.10
OpenAI-o1 52.78 65.45 56.28 60.00 41.96
Gemini-2.5-Pro 61.69 69.13 63.75 80.67 66.08
QwenVL2.5-32B 26.00 30.86 27.35 63.33 48.25
QwenVL2.5-72B 27.10 31.58 28.35 70.00 50.35
InternVL3.0-38B 26.90 29.24 27.55 68.00 45.10
InternVL3.0-78B 28.80 35.92 30.20 71.33 51.40
Baselines
QwenVL2.5-7B 19.99 22.56 20.70 55.33 28.47
InternVL3.0-8B 21.09 23.83 21.85 61.33 36.71
Advancing
     + GMAI-VL-R1 SFT 23.55 + 2.85 56.00 + 0.67 32.99 + 4.52
     + GMAI-VL-R1 RLT 23.80 + 3.10 57.33 + 2.00 34.03 + 5.56
     + Stage-I (ours) 24.48 26.90 25.15 + 4.45 62.00 + 6.67 36.36 + 7.89
     + Stage-II (ours) 25.80 28.52 26.55 + 5.85 66.00 + 10.67 38.81 + 10.34
     + Stage-I (ours) 26.14 28.52 26.80 + 4.95 69.33 + 8.00 43.36 + 6.65
     + Stage-II (ours) 25.93 31.05 27.35 + 5.50 70.00 + 8.67 48.95 + 12.24
Comparison with State-of-the Art Models. Table 1 also presents a comparison with leading models. Among open-source models, those enhanced with MedE2 demonstrate competitive or even superior performance despite having fewer parameters. For instance, on the MMMU-Health benchmark, QwenVL2.5-7B with MedE2 achieves an accuracy of 66%, outperforming its larger counterpart, QwenVL2.5-32B. The performance of InternVL3.0-8B with MedE2 has already surpassed InternVL3.0-38B and is slightly lower than InternVL3.0-78B. Similarly, on the MedXpertQA-MM and MMMU-Pro-Health benchmarks, the smaller models with MedE2 exhibit performance comparable to that of larger-scale models.
4.4 Ablation Studies
Larger Models, Greater Improvements Based on the preceding results, we observe that models exhibiting stronger initial performance tend to benefit more from reasoning elicitation. We hypothesize that larger models equipped with more extensive pretrained knowledge can derive increased benefits from our proposed MedE2 framework. To rigorously verify this hypothesis, we apply our training pipeline to QwenVL2.5-32B and QwenVL2.5-72B and evaluate them on MedXpertQA-MM. Notably, these two larger models are fine-tuned using LoRA. We select MedXpertQA-MM for evaluation due to its increased challenge and minimal data leakage risk, as it was carefully constructed using difficulty-based filtering and data augmentation. Figure 3 illustrates the accuracy variations across different model sizes at various stages. Following Stage-I, QwenVL2.5-72B demonstrates the most substantial performance gain, outperforming both QwenVL2.5-32B (3.65%) and QwenVL2.5-7B (4.45%). These findings are consistent with our observations in Section 3.2: models with more extensive prerequisite knowledge encoded within their parameter space are more capable of learning reasoning patterns from limited yet precisely structured exemplars. Combining the results shown in Table 1, we surprisingly find that QwenVL2.5-72B enhanced with MedE2 even surpasses QvQ-Max (40.45% versus 36.25%) on the MedXpertQA-MM benchmark, which is currently one of the most powerful models within the Qwen series. Although larger proprietary models such as o1 and Gemini-2.5-Pro still maintain an advantage on challenging benchmarks, the application of MedE2 significantly narrows the performance gap between open-source and proprietary models. These results further demonstrate the effectiveness of MedE2 in advancing multimodal reasoning capabilities for medical tasks. Detailed experimental results can be found in Appendix B.
Figure 3: The performance of QwenVL2.5-7B, QwenVL2.5-32B, and QwenVL2.5-72B on the MedXpertQA-MM benchmark. As the model size increases, the models demonstrate progressively greater benefits from the MedE2.
Text-only vs. Multimodal Elicitation. Previous studies have attempted to utilize multimodal training samples to activate multimodal reasoning capabilities. However, models optimized with such data often suffer from performance degradation on general vision tasks and impairment in linguistic abilities. Specifically, we compare text-only SFT, multimodal SFT, and their combination within Stage-I. In addition to evaluating on multimodal benchmarks (MedXpertQA-MM and MMMU-Health), we also assess performance on two language-focused benchmarks: MedQA and Medbullets. As illustrated in Figure 4, both text-only and multimodal SFT successfully elicit reasoning behaviors and improve overall performance. However, multimodal SFT yields relatively smaller improvements compared to text-only SFT. Notably, combining text-only and multimodal SFT results in reduced performance compared to the text-only SFT setting, particularly evident in the text tasks. This finding confirms that eliciting reasoning abilities through multimodal data may compromise the original language capabilities of the model.
Figure 4: Comparison of performance on Med-XpertQA-MM using various strategies for eliciting reasoning behaviors.
Figure 5: The Accuracy of models with MedE2 on MedXpertQA-MM (upper part) and MMMU-Pro-Health (lower part) with inference-time scaling.
Benifition when Applying Inference Time Scaling. To investigate the effectiveness of MedE2, in scaling up inference time, we report the results of majority-vote accuracy and pass@n accuracy on MedXpertQA-MM and MMMU-Pro-Health across 32 computational runs with temperature =0.5. The results, illustrated in Figure 5, compare the performance of three models under Pass@n and majority voting (MV@n). Since both datasets are in a multiple-choice format with a limited answer space, model performance consistently improves as the number of attempts increases. Although the performance gain from majority voting on the MedXpertQA-MM dataset is modest due to its inherent difficulty, models trained with MedE2 still demonstrate a significant and stable advantage as the number of computational attempts increases. On MMMU-Pro-Health, employing majority voting yields notable performance enhancements. Interestingly, the accuracy of QwenVL2.5-7B initially rises to a peak but then slightly declines as computational runs increase. In contrast, models trained with MedE2 exhibit a steady performance improvement. These findings suggest that MedE2 enhances model output consistency, particularly evident during majority voting.
Figure 6: Case illustration of models’ reasoning capability in tackling clinical decision task.
4.5 Qualitative Analysis
Figure 6 demonstrates the reasoning capabilities of QwenVL2.5-72B+Stage2 through its coherent analysis of determining appropriate follow-up examinations based on the patient’s current clinical presentation. In real-world clinical scenatios, making an accurate diagnosis requires clinicians to integrate findings from multiple examinations to progressively narrow the differential diagnosis. Therefore, formulating an efficient and precise examination plan tailored to the patient’s chief complaint is critical. Taking a closer look to the case, a notable observation is that the reasoning process is not constrained by the given options. Instead, it reflects a stepwise, hypothesis-driven diagnostic approach rather than merely discussing each option sequentially. Specifically, the model first identifies potential diagnoses (e.g., Thrombocytopenia-Absent Radius syndrome and Fanconi anemia) and then systematically excludes one by evaluating evidence from the radiographic findings. Decisions regarding subsequent examinations are logically derived from the provisional diagnosis, aligning closely with realistic clinical workflows. This observed capability highlights the potential of MedE2 to serve as an effective post-trainig recipe to narrow the gap between current model performance and practical clinical application. We also remove the multiple-choice options and allow the model to respond freely, which can be seen in the Figure 7 of Appendix C.
5 Conclusion and Discussion
Conclusion. In this study, we propose a novel reinforcement-learning-free training pipeline, MedE2, designed to progressively elicit and enhance multimodal reasoning capabilities in medical domains. MedE2 comprises two distinct stages: it first performs supervised fine-tuning on carefully curated text-only medical reasoning data, followed by preference-based optimization using multimodal data through Direct Preference Optimization. To support medical reasoning training, we constructed a dataset of 3.5K questions spanning both textual and multimodal clinical scenarios. Extensive experiments demonstrate the effectiveness and reliability of MedE2 in advancing the reasoning abilities of medical multimodal models. The demonstrated scalability and consistent improvements under inference-time scaling highlight its broad applicability across different model architectures and parameter scales. We hope that this early exploration into multimodal reasoning for medicine will inspire further research into specialized clinical reasoning capabilities.
Limitation and Impact. We hope that our post-training strategy can open new avenues for developing multimodal reasoning models in the medical domain. Despite the promising performance improvements demonstrated by our pipeline on current models, the lack of multimodal large language models specifically tailored to the medical domain presents significant challenges. Moreover, although current evaluation benchmarks are designed to resemble clinical scenarios, a notable gap remains. For example, many benchmark cases are crafted with sufficient cues that allow a medical specialist to deduce the correct diagnosis. In contrast, it would be challenging to create such a concise, complete and coherent case report in real clinical settings. Therefore, collecting real-world data and understanding how models generalize their reasoning to more complex and challenging clinical contexts remains a critical direction for future research.
Appendix A Prompt Templates and Data Formats Adopted in MedE2
[…]
Appendix B Experimental Results with Different Model Scales on MedXpertQA-MM
Our method demonstrates more significant performance on larger-scale models, as shown in Table 2. These findings are consistent with our observations in Section 3.2: models with more extensive prerequisite knowledge encoded within their parameter space are more capable of learning reasoning patterns from limited yet precisely structured exemplars.
Table 2: Performance comparison of QwenVL2.5-32B and QwenVL2.5-72B with MedE2 across different task categories on the MedXpertQA-MM benchmark. As shown in the figure, the accuracy of each subtask improves, with the larger 72B model exhibiting a more pronounced increase compared to the 32B model.
Method Treatment Basic Science Diagnosis Reasoning Understanding Overall $\Delta$
QwenVL2.5-32B 28.13 25.78 27.61 26.07 30.87 27.40 +0.00
   +Stage-I (ours) 32.37 33.14 29.94 30.15 33.39 31.05 +3.65
   +Stage-II (ours) 35.92 34.15 30.85 31.62 34.88 32.50 +5.10
QwenVL2.5-72B 30.36 24.08 28.86 27.11 31.59 28.35 +0.00
   +Stage-I (ours) 40.40 41.93 38.62 39.21 40.61 39.60 +11.25
   +Stage-II (ours) 41.52 42.49 39.45 39.63 42.60 40.45 +12.10
Appendix C Case Study
Figure 7: Case illustration of models’ reasoning capability in the open-ended task.
Figure 8: Case illustration of the differences in the responses of models in Stage-I and Stage-II. Blue text in the figure indicates visual hallucinations generated by the models.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract clearly states our main contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include the limitations of our work in Conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theoretical paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We explain our settings in Experiments and hyperparameters in Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Anonymous URL offered.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details are summarized in Appendix A, with explanations on data split in Experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We show the standard error in most training curves, with an average of over three random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We explain it on Experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read and understood the code of ethics; and have done our best to conform.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work only involves public dataset. It does no impact the society at large, beyond improving our utilization on medical CT images.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Since we train models on public CT datasets, our work poses no risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We are not shipping our code with any source code or binary files from any other existing libraries, so there are no concerns over getting permission or including a license. We did cite open-sourced libraries in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will release our code base with included readme files. We do not ship any source code or binary files from any other existing libraries.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM does not impact the core methodology, scientific rigorousness, or originality of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.